# TCP-Diffusion: A Multi-modal Diffusion Model for Global Tropical Cyclone Precipitation Forecasting with Change Awareness

### v1

@@ Abstract
Precipitation from tropical cyclones (TCs) can cause disasters such as flooding, mudslides, and landslides. Predicting such precipitation in advance is crucial, giving people time to prepare and defend against these precipitation-induced disasters. Developing deep learning (DL) rainfall prediction methods offers a new way to predict potential disasters. However, one problem is that most existing methods suffer from cumulative errors and lack physical consistency. Second, these methods overlook the importance of meteorological factors in TC rainfall and their integration with the numerical weather prediction (NWP) model. Therefore, we propose Tropical Cyclone Precipitation Diffusion (TCP-Diffusion), a multi-modal model for global tropical cyclone precipitation forecasting. It forecasts TC rainfall around the TC center for the next 12 hours at 3 hourly resolution based on past rainfall observations and multi-modal environmental variables. Adjacent residual prediction (ARP) changes the training target from the absolute rainfall value to the rainfall trend and gives our model the capability of rainfall change awareness, reducing cumulative errors and ensuring physical consistency. Considering the influence of TC-related meteorological factors and the useful information from NWP model forecasts, we propose a multi-model framework with specialized encoders to extract richer information from environmental variables and results provided by NWP models. The results of extensive experiments show that our method outperforms other DL methods and the NWP method from the European Centre for Medium-Range Weather Forecasts (ECMWF).
Introduction
@@
Tropical cyclone (TC) precipitation refers to the substantial rainfall events that accompany the formation, development, and movement of tropical cyclones. These precipitation events can lead to severe disasters in the affected regions, like heavy rainfall, flooding, mudslides, and landslides. These TC rainfall-related disasters cause more economic losses and deaths than strong winds (Wendler-Bosco and Nicholson 2022), yet most TC forecasting efforts focus only on track and wind intensity. Therefore, it is important to accurately predict the precipitation around the TC center.
@@
Figure 1: The TC precipitation forecasting task. The red box represents historical data, which is considered to be the input of the forecasting model. The blue box represents the TC rainfall that should be predicted. This paper just focuses on the prediction of TC rainfall.
@@
Numerical weather prediction (NWP) is widely used for various weather forecasting tasks, including TC forecasting. It simulates atmospheric dynamics, thermodynamics, and physical processes by solving the physical and mathematical equations that describe changes in wind, temperature, humidity, and pressure (Bai et al. 2020). It has achieved impressive results, but the development in NWP research has been gradual (Benjamin et al. 2018). Additionally, not all physical mechanisms of TCs are fully understood, which prevents the formulation of perfect physical and mathematical equations for NWP (Xu et al. 2024). Furthermore, NWP typically runs on supercomputer platforms, consuming extensive computational resources and time for predictions, when predicting a high-resolution result. Nonlinear simulation also poses a significant challenge for NWP. Fortunately, the development of deep learning (DL) offers solutions to these NWP challenges. DL-based methods excel at modeling nonlinear processes, and the cost of model training and inference is relatively low, requiring only a few graphics processing units (GPUs) for regular AI tasks.
@@
DL methods have been applied to many meteorological tasks, like TC track and intensity prediction (Huang et al. 2022, 2023), TC rainfall downscaling (Vosper et al. 2023), global meteorological variable prediction (Bi et al. 2023; Lam et al. 2023), and precipitation nowcasting (Bai et al. 2022). For precipitation nowcasting, several works have used DL since 2013. Initially, researchers typically used a U-shape architecture network to predict rainfall (Gao et al. 2022; Xu et al. 2021). This was an interesting and meaningful attempt in the precipitation nowcasting task. However, due to the blurry predictions from the U-shape architecture network, researchers began using the generative adversarial network (GAN) model to obtain more realistic rainfall predictions (Tian et al. 2019). Recently, due to issues with mode collapse and training instabilities in GAN models (Gao et al. 2024), researchers have applied diffusion models (DMs) for better rainfall prediction (Asperti et al. 2023).
However, the TC precipitation forecasting task does not attract much attention from researchers, which is also important for people to defend against related disasters. Additionally, there are differences between TC precipitation forecasting and regular precipitation forecasting. For example, regular precipitation forecasting usually focuses on a specific fixed region, where the area being predicted does not move with the rainfall. In contrast, in TC precipitation forecasting, the primary region of interest moves with the TC. Therefore, it is inappropriate to directly apply methods designed for regular precipitation forecasting to TC rainfall prediction. A framework specifically designed for TC precipitation is necessary.
Three aspects could be considered: changing the training goal, extracting richer meteorological information from related TC data, and integrating it with NWP.
@@
Changing the training goal: most precipitation forecasting methods predict the absolute value of rainfall. However, future rainfall can be understood as the sum of the current rainfall and change in rainfall ($\Delta$ Rainfall) over time. Specifically:
$Rainfall_{Future}=Rainfall_{Current}+\Delta Rainfall$ (1)
@@
We can train a DL model to predict $\Delta$ Rainfall and use Equation 1 to obtain the absolute value of future rainfall, rather than predicting the absolute value of future rainfall directly. This mechanism is advantageous because predicting the change in TC rainfall helps reduce cumulative errors and ensures physical consistency. We define this capability of a model as change awareness. Similar mechanisms are used in some NWP methods to improve the accuracy and stability of meteorological forecasts (Kalnay 2003). Therefore, it is worth considering this mechanism in the TC rainfall prediction task.
@@
Extracting richer meteorological information: most previous DL precipitation forecasting methods attempt to extract sufficient information solely from the rainfall value data. However, the information in rainfall data alone is not adequate for DL models to learn the patterns of TC rainfall. Various TC-related data can help people and DL models understand and predict the tendency of TC rainfall, shown in the red box in Figure 1. For example, rainfall is driven by wind, so the distribution of TC wind is important for TC rainfall prediction. Therefore, in addition to directly extracting information from rainfall data, it is also important to focus on the information derived from the relationships between meteorological factors.
Integrating with NWP models: researchers are trending towards building a suitable “bridge” between DL models and NWP models. Research on this “bridge” can be broadly divided into two categories: 1. using DL methods to replace a part of or all NWP processes (Bi et al. 2023), and 2. focusing on correcting the predictions provided by NWP models (Harris et al. 2022; Zhu et al. 2022). The first category uses the power of DL models’ nonlinear modeling to replace the complex processes of NWP models, making the prediction process easier and more effective. Methods in this category can learn from historical data, such as TC rainfall data from the past few hours. Models in the second category can more directly extract useful information from prediction results calculated by physical and mathematical equations. Therefore, considering the advantages of both categories, it is beneficial to build a framework that can extract information from various historical variables while also using specific information from NWP results to guide better predictions (Antonio et al. 2024).
@@
Based on these above potentially useful ideas, we propose TCP-Diffusion, a multi-modal diffusion model for global tropical cyclone precipitation forecasting with change awareness. The contributions are summarized as follows:
@@
  • Adjacent Residual Prediction (ARP) is proposed to make our model focus on the rainfall change between adjacent time steps and have the capability of rainfall change awareness. This mechanism can reduce cumulative errors and ensure physical consistency, making the results more realistic and accurate.
@@
Figure 2: The framework of TCP-Diffusion. U-shape architecture (in the middle) includes the U-Net encoder and U-Net decoder. In inference stage (black+red arrows), the “N Steps” (at the top) means our model will use N loop denoising processes to obtain the result $\hat{D}_{y}$ and perform Post Process (Equation 1) to get the final predicted TC rainfall $\hat{Y}$. During the training stage (black+green arrows), “N Steps” and Post Process will not be used. The Output will be used to calculate the loss and update the parameters of TCP-Diffusion.
@@
Different from regular precipitation forecasting, where the predicted regions are fixed, TC precipitation forecasting focuses on the $10^{\circ}$ by $10^{\circ}$ rainfall field around the TC center, which moves with the TC. We denote the final output of our work as $\hat{Y}=\{\hat{y}_{n+1},\hat{y}_{n+2},\ldots,\hat{y}_{n+t},\ldots,\hat{y}_{n+m}\}$, where $\hat{y}_{n+t}$ represents the rainfall prediction around the TC center $t$ time steps into the future. Here, $n$ represents the input time steps and $m$ represents the output time steps. The input data we used can be divided into two parts, as shown in the left part of Figure 2: the first part is historical observation data ($X_{historical}=\{X_{1}^{h},X_{2}^{h},\ldots,X_{t}^{h},\ldots,X_{n}^{h}\}$), where $X_{t}^{h}$ is the input data at the $t$ time step and it includes the rainfall value data, $\Delta$ Rainfall, the 2-Dimension environment data, and scalar TC variables, like intensity. The second part is the future prediction data from an NWP system ($X_{future}=\{X_{1}^{f},X_{2}^{f},\ldots,X_{t}^{f},\ldots,X_{m}^{f}\}$), where $X_{t}^{f}$ denotes the prediction data at future $t$ time step, which is provided by ECMWF (ECMWF 2017). Overall, imagine a TC developing, just like shown in Figure 1. We obtain all the input data $X=\{X_{historical},X_{future}\}$ and then we input them to our model TCP-Diffusion. TCP-Diffusion will extract the TC information from the input data and make a reasonable and accurate rainfall prediction $\hat{Y}$. We expect $\hat{Y}$ could be as close as possible to the real TC rainfall ($Y=\{y_{n+1},y_{n+2},\ldots,y_{n+t},\ldots,y_{n+m}\}$) as much as possible, both in the value of the rainfall and the visual authenticity, where $y_{n+t}$ represents the real rainfall around the TC center at the future $t$ time step.
Data
We collected TC various rainfall-related data for our TCP-Diffusion model to comprehensively represent TC rainfall. A total of 1877 TCs spanning from 1980 to 2020 are collected, covering the six major ocean areas. These TC data are divided into three sets: training set (1751 TCs), validation set (87 TCs), and test set (126 TCs from 2018 to 2020). We divided the data into two parts: $X_{historical}$ and $X_{future}$, which are shown as follows. More details of the data we used could be checked in the Supplementary Material and we will also release all these data after that the paper is accepted.
Historical Data
Rainfall Data.
It is the main data of this task, denoted as $X_{rain}$. It is collected from Multi-Source Weighted-Ensemble Precipitation (MSWEP) (Beck et al. 2019), a global precipitation product with a 3‑hourly ${0.1}^{\circ}$ resolution available from 1979 to about 3 hours from real-time. We focus only on the rainfall field around the TC center, so we crop and obtain rectangular data covering a $10^{\circ}$ by $10^{\circ}$ region around the TC center.
Environment Data.
Rainfall data provide direct information related to this task, but environmental factors should not be neglected when modeling TC rainfall. Thus, we also collect critical meteorological variables including 2D environment data (ERA5 surface data and ERA5 pressure level data) and Scalar TC variables. The surface data, denoted as $X_{SfEnv}$, like 2m temperature. ERA5 pressure level data, denoted as $X_{PlEnv}$, like specific humidity. $X_{SfEnv}$ and $X_{PlEnv}$ are obtained from ERA5 (Hersbach et al. 2020). We perform the same operations as with $X_{rain}$, focusing on the $10^{\circ}$ by $10^{\circ}$ region around the TC center. Apart from these above normal environment data, Scalar TC variables are quite related to the TC rain. We consider it as internal environment data, which is denoted as $X_{InEnv}$. It includes TC intensity, movement velocity, month and the location. $X_{InEnv}$ is collected and calculated from the International Best Track Archive for Climate Stewardship (IBTrACS) resource (Knapp et al. 2010).
Future Prediction Data.
The future prediction data generated by NWP contain explicit physical laws, which are precisely what deep learning methods lack. However, DL models are good at extracting information from data. Therefore, we also consider $X_{future}$ as input data for our model, providing future information and physical laws to some extent. All $X_{future}$ is provided by ERA5-IFS forecasts (ECMWF 2017). There are many variables at the 200 or 850 hPa pressure levels, such as the U-component of wind and the V-component of wind.
Method
@@
TCP-Diffusion is a temporal DM-based method. We call the main component Environment Awareness 3DUNet (EA-3DUNet) (in section EA-3DUNet). As shown in Figure 2, it can use Historical $Data_{2d}$ Encoder, Historical $Data_{1d}$ Encoders and Future $Data_{2d}$ Encoder to extract TC rainfall information from rainfall related data, 2D environmental factor data, scalar TC variable and future prediction data. 3DUNet is designed to further extract spatiotemporal future from the information provided by different encoders. The inherent power of DMs ensures that the prediction resembles real rainfall data, both in field shape and detail. The information from all the inputs helps the predicted rainfall values closely match the real rainfall data. In the following section, we will introduce the details of TCP-Diffusion. More details of the method could be checked in the Supplementary Material and we will also release all these Python codes after that the paper is accepted.
Adjacent Residual Prediction (ARP)
@@
In the TC rainfall forecasting task, predicting the rainfall value directly is difficult because weather systems exhibit chaotic characteristics. Some NWP methods improve the accuracy and stability of forecasts by keeping the weather change as their prediction goal (Kalnay 2003). Inspired by this idea, we find the capability of change awareness is helpful to our model, so we replace the direct prediction of rainfall value with predicting the adjacent residual $\Delta_{x}^{t}$ between rainfall at adjacent time steps ($\Delta$ Rainfall). We call this mechanism Adjacent Residual Prediction. $\Delta_{x}^{t}$ can be calculated as:
@@
where $X_{rain}^{t}$ and $X_{rain}^{t-1}$ are the original rainfall data at $t$ time step and $t-1$ time step respectively. In our work, apart from the absolute rainfall values data, we add adjacent residual sequence $D_{x}=\{\Delta_{x}^{1},\Delta_{x}^{2},\ldots,\Delta_{x}^{t},\ldots,\Delta_{x}^{n}\}$, where $\Delta_{x}^{t}$ is calculated by the Equation 2, which shown as the $\Delta$ Rainfall in Figure 2. Additionally, the direct output sequence of our DL model is $\hat{D}_{y}=\{\hat{\Delta}_{y}^{n+1},\hat{\Delta}_{y}^{n+2},\ldots,\hat{\Delta}_{y}^{n+t},\ldots,\hat{\Delta}_{y}^{n+m}\}$, where $\hat{\Delta}_{y}^{n+t}$ is the adjacent residual we predict between rainfall of time step $n+t-1$ and $n+t$. The final prediction $\hat{Y}=\{\hat{y}_{n+1},\hat{y}_{n+2},\ldots,\hat{y}_{n+t},\ldots,\hat{y}_{n+m}\}$ can be computed as:
@@
$\hat{Y}$ is obtained by the accumulation of the latest historical TC rainfall data $X_{rain}^{n}$ and the adjacent residual prediction $\hat{\Delta}_{y}$.
Diffusion Model
The prediction process of DMs involves removing noise from a completely random noise $r$ step by step, shown in Figure 2’s $N$ Steps. Based on the DMs’ generation process, our training goal is to teach our model to perform accurate denoising based on the specific step $s$ and the given information. Specifically, the training goal is to make the noise $\hat{r}_{s}$ predicted by our model as close as possible to the noise $r_{s}$ added to the target at the $s$-th step, $s\in\{1,2,\ldots,N\}$. $N$ is a hyper-parameter.
DMs typically include two processes: a forward noising process and a reverse denoising process. The forward process has no trainable parameters, so it does not require training a DL model. According to the training goal mentioned above, we need to obtain the sequence $D_{y}^{s}$, where $D_{y}^{s}$ is $s$-th-step noised real adjacent residual $D_{y}=\{\Delta_{y}^{n+1},\Delta_{y}^{n+2},\ldots,\Delta_{y}^{n+t},\ldots,\Delta_{y}^{n+m}\}$. Follow (Ho et al. 2022), $D_{y}^{s}$ could be calculated using a random noise $r_{s}\sim\mathcal{N}(0,1)$, step $s$, and the $D_{y}^{0}$ (the original data without noise). The forward process is defined as follows:

### v2

@@ Abstract
Deep learning methods have made significant progress in regular rainfall forecasting, yet the more hazardous tropical cyclone (TC) rainfall has not received the same attention. While regular rainfall models can offer valuable insights for designing TC rainfall forecasting models, most existing methods suffer from cumulative errors and lack physical consistency. Additionally, these methods overlook the importance of meteorological factors in TC rainfall and their integration with the numerical weather prediction (NWP) model. To address these issues, we propose Tropical Cyclone Precipitation Diffusion (TCP-Diffusion), a multi-modal model for forecasting TC precipitation given an existing TC in any location globally. It forecasts rainfall around the TC center for the next 12 hours at 3 hourly resolutions based on past rainfall observations and multi-modal environmental variables. Adjacent residual prediction (ARP) changes the training target from the absolute rainfall value to the rainfall trend and gives our model the capability of rainfall change awareness, reducing cumulative errors and ensuring physical consistency. Considering the influence of TC-related meteorological factors and the useful information from NWP model forecasts, we propose a multi-model framework with specialized encoders to extract richer information from environmental variables and results provided by NWP models. The results of extensive experiments show that our method outperforms other DL methods and the NWP method from the European Centre for Medium-Range Weather Forecasts (ECMWF).
Machine Learning, ICML
1 Introduction
@@
Tropical cyclone (TC) precipitation refers to substantial rainfall events that accompany the formation, development, and movement of TCs. These precipitation events can lead to severe disasters in the affected regions, like flooding, mudslides, and landslides. These TC rainfall-related disasters cause more economic losses and deaths than strong winds (Wendler-Bosco & Nicholson, 2022), yet most TC forecasting research focuses only on TC track and intensity while neglecting the TC precipitation. If the TC precipitation can be accurately predicted, preemptive measures may be taken to mitigate the disasters brought by TC rainfall, protecting people’s lives and property. Therefore, it is important to accurately predict the precipitation around the TC center.
@@
Figure 1: The difference between the regular precipitation forecasting (a) and TC (b) precipitation forecasting. Regular precipitation forecasting (a) usually focuses on a specific fixed region, where the predicted area (blue box) does not move with the rainfall area (black dashed boxes). In TC precipitation forecasting (b), the primary region of interest (blue boxes) moves with the TC. Sub-figure (c) represents the main prediction process of our model. The light blue box represents historical data, which is considered to be the input of the forecasting model. The light orange box represents the TC rainfall that should be predicted. This paper just focuses on the prediction of TC rainfall.
@@
Numerical weather prediction (NWP) is widely used for various weather forecasting tasks, including TC rainfall forecasting. It simulates atmospheric dynamics, thermodynamics, and physical processes by solving the physical and mathematical equations that describe changes in wind, temperature, humidity, and pressure (Bai et al., 2020). It has achieved impressive results, but the development in NWP research has been gradual (Benjamin et al., 2018). Additionally, not all physical mechanisms of TCs are fully understood, which prevents the formulation of perfect physical and mathematical equations for NWP (Xu et al., 2024). Furthermore, NWP typically runs on supercomputer platforms, consuming extensive computational resources and time for predictions, when predicting a high-resolution result. Nonlinear simulation also poses a significant challenge for NWP. Fortunately, the development of deep learning (DL) offers solutions to these NWP challenges. DL-based methods excel at modeling nonlinear processes, and the cost of model training and inference is relatively low, requiring only a few Graphics Processing Units for regular AI tasks.
@@
DL methods have been applied to many meteorological tasks, like TC track and intensity prediction (Huang et al., 2022, 2023; Zhang et al., 2025), TC rainfall downscaling (Vosper et al., 2023), global meteorological variable prediction (Bi et al., 2023; Lam et al., 2023), and precipitation nowcasting (Bai et al., 2022; Yang et al., 2024). For regular precipitation nowcasting, several works have used DL. Some previous works used a U-Net architecture network to predict rainfall (Gao et al., 2022; Xu et al., 2021). This could perform well on metrics like mean squared error, but gave predictions that were spatially smoother than real TC rainfall. Some research used the generative adversarial network (GAN) model to obtain more realistic rainfall predictions (Tian et al., 2019). Recently, due to issues with mode collapse and training instabilities in GAN models (Gao et al., 2024), researchers have applied diffusion models (DMs) for better rainfall prediction (Asperti et al., 2023; Addison et al., 2022, 2024).
These previous precipitation forecast models provide a good reference for constructing TC precipitation models. However, there are notable differences between the TC precipitation forecasting task that we consider here and regular precipitation forecasting, as shown in Figure 1.(a) and (b). For example, regular precipitation forecasting usually focuses on a specific fixed region, where the area being predicted does not move with the rainfall. In contrast, in this work, we focus on forecasting the TC precipitation, where the primary region of interest and the corresponding environmental factors, such as terrain, will change dynamically with the movement of the TC. Therefore, it is valuable to investigate whether better methods can be developed for this task. We consider the three following approaches to improving TC rainfall prediction skills:
@@
Changing the training goal: most precipitation forecasting methods predict the absolute value of rainfall. However, future rainfall can be understood as the sum of the current rainfall and change in rainfall ($\Delta Rainfall_{Future}$) over time, which is shown in Figure 2.(c). Specifically:
$Rainfall_{Future}=Rainfall_{Current}+\Delta Rainfall_{Future}$ (1)
@@
We can train a DL model to predict $\Delta$ Rainfall and use Equation 1 to obtain the absolute value of future rainfall, rather than predicting the absolute value of future rainfall directly. This mechanism is advantageous because predicting the change in TC rainfall not only helps reduce cumulative errors but also ensures physical consistency–meaning that changes in rainfall intensity and spatial patterns should align to some extent with historically observed trends. We define this capability of a model as change awareness. Similar mechanisms are used in some NWP methods (Kalnay, 2003) and global weather forecasting models (Price et al., 2025; Nguyen et al., 2024; Oskarsson et al., 2024; Hu et al., 2023) to improve the accuracy and stability of meteorological forecasts. Therefore, it is worth considering this mechanism in the TC rainfall prediction task.
@@
Extracting richer meteorological information: most previous DL precipitation forecasting methods attempt to extract sufficient information solely from the rainfall value data. However, the information in rainfall data alone is not adequate for DL models to learn the patterns of TC rainfall. Various TC-related data can help people and DL models understand and predict the tendency of TC rainfall, shown in the light blue box in Figure 1.(c). For instance, since rainfall is wind-driven, TC wind distribution is critical. Thus, beyond direct rainfall data extraction, modeling synergistic effects of meteorological factors (e.g., wind-rainfall dynamics) is also essential to capture TC rainfall mechanisms.
Integrating with NWP models: DL models can make use of global NWP predictions to enhance forecasts of specific phenomena (Harris et al., 2022; Zhu et al., 2022; Antonio et al., 2024). If the NWP forecasts are produced using physical equations, such postprocessing approaches allow useful information to be extracted from the embedded understanding. Therefore, we evaluate using specific information from NWP results to guide better predictions.
@@
Based on these above potentially useful ideas, we propose TCP-Diffusion, a multi-modal diffusion model for global tropical cyclone precipitation forecasting with change awareness. This model can simultaneously forecast TC rainfall for multiple future time steps, at 3-hourly intervals, within the next 12 hours. The contributions are summarized as follows:
@@
  • Adjacent Residual Prediction (ARP) is proposed to make our model focus on the rainfall change between adjacent time steps, learn a coherent temporal evolution of the TC rainfall, and have the capability of rainfall change awareness. This mechanism can reduce cumulative errors and ensure physical consistency, making the results more realistic and accurate.
@@ 2 Task Definition
Figure 2: The framework of TCP-Diffusion. Sub-figure (a) shows the inference stage of TCP-Diffusion (black+red arrows). Our model will use $N$ loop denoising processes to obtain the future rainfall change $\Delta Rainfall_{Future}$ from a noise input. The sub-figure (b) shows the model structure of the denoising process and also shows the training stage of TCP-Diffusion (black+green arrows). In the training stage, the Output will be used to calculate the loss and update the parameters of TCP-Diffusion (green arrow). In the inference stage, the Output will be used to get the future rainfall change $\Delta Rainfall_{Future}$ with noise at Step $s-1$ (red arrow). The sub-figure (c) shows the process (Equation 1) of getting the final predicted TC rainfall $Rainfall_{Future}$ with the latest rainfall $Rainfall_{Current}$ and the future rainfall change $\Delta Rainfall_{Future}$.
@@
Different from regular precipitation forecasting, where the predicted regions are fixed, our TC precipitation forecasting task focuses on the $10^{\circ}$ by $10^{\circ}$ rainfall field around the TC center, which moves with the TC. We denote the final output of our work as $\hat{Y}=\{\hat{y}_{n+1},\hat{y}_{n+2},\ldots,\hat{y}_{n+t},\ldots,\hat{y}_{n+m}\}$, where $\hat{y}_{n+t}$ represents the rainfall prediction around the TC center $t$ time steps into the future. Here, $n$ represents the input time steps and $m$ represents the output time steps. The input data we used can be divided into two parts, as shown in the left part of Figure 2.(b): the first part is historical observation data ($X_{historical}=\{X_{1}^{h},X_{2}^{h},\ldots,X_{t}^{h},\ldots,X_{n}^{h}\}$), where $X_{t}^{h}$ is the input data at the $t$ time step and it includes the rainfall value data, $\Delta$ Rainfall, the 2-Dimension environment data, and scalar TC variables, like TC intensity and track. The second part is the future prediction data from an NWP system ($X_{future}=\{X_{1}^{f},X_{2}^{f},\ldots,X_{t}^{f},\ldots,X_{m}^{f}\}$), where $X_{t}^{f}$ denotes the prediction data at future $t$ time step. Overall, imagine a TC developing, just like shown in Figure 1.(c). We obtain all the input data $X=\{X_{historical},X_{future}\}$ and then we input them to our model TCP-Diffusion. TCP-Diffusion will extract the TC information from the input data and make a rainfall prediction $\hat{Y}$. The aim is for $\hat{Y}$ to be as close as possible to the real TC rainfall ($Y=\{y_{n+1},y_{n+2},\ldots,y_{n+t},\ldots,y_{n+m}\}$, where $y_{n+t}$ represents the real rainfall around the TC center at the future $t$ time step), both in terms of the overall intensity and the spatial and temporal structure.
3 Model Design
@@
TCP-Diffusion is a temporal DM-based method. The denoising process is performed by a neural network, which we call Environmentally-Aware 3DUNet (EA-3DUNet) (described in detail below). As shown in Figure 2.(b), it can use Historical $Data_{2d}$ Encoder, Historical $Data_{1d}$ Encoders and Future $Data_{2d}$ Encoder to extract TC rainfall information from rainfall related data, 2D environmental factor data, scalar TC variable and future prediction data. 3DUNet is designed to further extract spatiotemporal future features from the information provided by different encoders. The inherent abilities of DMs enables making predictions that have structure that resembles real rainfall data. The information from all the inputs helps the predicted rainfall values more closely match the real rainfall data. In the following section, we will introduce the details of TCP-Diffusion. We have released the code at https://github.com/Zjut-MultimediaPlus/TCP-Diffusion.
3.1 Adjacent Residual Prediction (ARP)
@@
In the TC rainfall forecasting task, predicting the rainfall value directly is difficult because weather systems exhibit chaotic characteristics. Some NWP methods improve the accuracy and stability of forecasts by keeping the weather change as their prediction goal (Kalnay, 2003). Inspired by this idea, we find the capability of change awareness is helpful to our model, so we replace the direct prediction of rainfall value with predicting the adjacent residual $\Delta_{x}^{t}$ between rainfall at adjacent time steps ($\Delta$ Rainfall). We call this mechanism Adjacent Residual Prediction. $\Delta_{x}^{t}$ can be calculated as:
@@
where $X_{rain}^{t}$ and $X_{rain}^{t-1}$ are the absolute rainfall data at $t$ time step and $t-1$ time step respectively. In our work, in addition to the absolute rainfall data, we add adjacent residual sequence $D_{x}=\{\Delta_{x}^{1},\Delta_{x}^{2},\ldots,\Delta_{x}^{t},\ldots,\Delta_{x}^{n}\}$, where $\Delta_{x}^{t}$ is calculated by the Equation 2, which shown as the $\Delta$ Rainfall in Figure 2.(b). The direct output sequence of our DL model is $\hat{D}_{y}=\{\hat{\Delta}_{y}^{n+1},\hat{\Delta}_{y}^{n+2},\ldots,\hat{\Delta}_{y}^{n+t},\ldots,\hat{\Delta}_{y}^{n+m}\}$, where $\hat{\Delta}_{y}^{n+t}$ is the adjacent residual we predict between rainfall of time step $n+t-1$ and $n+t$. The final prediction $\hat{Y}=\{\hat{y}_{n+1},\hat{y}_{n+2},\ldots,\hat{y}_{n+t},\ldots,\hat{y}_{n+m}\}$ can be computed as:
@@
$\hat{Y}$ is obtained by the accumulation of the latest historical TC rainfall data $X_{rain}^{n}$ and the ARP $\hat{\Delta}_{y}$.
3.2 Diffusion Model
The prediction process of DMs involves removing noise from a completely random noise $r$ step by step, shown in Figure 2.(a). Based on the DMs’ generation process, our training goal is to teach our model to perform accurate denoising based on the specific step $s$ and the given information. Specifically, the training goal is to make the noise $\hat{r}_{s}$ predicted by our model as close as possible to the noise $r_{s}$ added to the target at the $s$-th step, $s\in\{1,2,\ldots,N\}$. $N$ is a hyper-parameter.
DMs typically include two processes: a forward noising process and a reverse denoising process. The forward process has no trainable parameters, so it does not require training a DL model. According to the training goal mentioned above, we need to obtain the sequence $D_{y}^{s}$, where $D_{y}^{s}$ is $s$-th-step noised real adjacent residual $D_{y}=\{\Delta_{y}^{n+1},\Delta_{y}^{n+2},\ldots,\Delta_{y}^{n+t},\ldots,\Delta_{y}^{n+m}\}$. Following (Ho et al., 2022), $D_{y}^{s}$ could be calculated using a random noise $r_{s}\sim\mathcal{N}(0,1)$, step $s$, and the $D_{y}^{0}$ (the original data without noise). The forward process is defined as follows: